OpenStudy (waheguru):

a hemisphere has a radius of 17 what is the surface area

OpenStudy (mertsj):

Surface area of a sphere = 4pi r^2 Plug in 17 in place of r and then divide the answer by 2 because it is a hemisphere.

OpenStudy (waheguru):

what is it i wanna see if my techer marked wroing

OpenStudy (mertsj):

What did you get?

OpenStudy (anonymous):

A=4*pi*r^2 A/2=2*pi*r^2=2*3.14*17*17=1816 plus pi*r^2

OpenStudy (waheguru):

i got 1816

OpenStudy (mertsj):

Yep. If you round it. actually I got 1815.84

OpenStudy (anonymous):

908+1816=2724

OpenStudy (mertsj):

Did you perhaps omit the units?

OpenStudy (waheguru):

how?

OpenStudy (mertsj):

Why are you adding cinar?

OpenStudy (waheguru):

i forgot the bottom surface

OpenStudy (anonymous):

cut a watermelon two pieces think about red area not only green area

OpenStudy (waheguru):

ahhhhhhhhhhhhhhhhhhhhhh

OpenStudy (waheguru):

thx

OpenStudy (anonymous):

yw